OpenStudy (anonymous):

HELP!!! a line has a slope of -2/3 and a y-intercept of (0,3).What is the value of y when x=4?

OpenStudy (anonymous):

HELP!!!

OpenStudy (anonymous):

\[y = mx + b \] where m is the slope and b is the y-intercept \[y = -2/3(4) + 3\]\[\rightarrow y = \frac{-8}{3} + \frac{9}{3}\] the 9/3 represents the 3 \[y = \frac{1}{3}\]

OpenStudy (anonymous):

do you just plug in a random number then simplify and work from there?

OpenStudy (anonymous):

you just plug in the x value. Because it's y = slope (x value) + y-intercept. You had the slope and the y-intercept, so you just had to plug in the x value and solve for y.

OpenStudy (anonymous):

ooooooooooohhhh i get it!! Thank you!!

OpenStudy (anonymous):

cool :) no problem